(-3 __ ) is a solution of y = 4 -3x.
Answer by jim_thompson5910(35256)

y = 4 -3x

y = 4 -3(-3)

y = 4 + 9

y = 13

So (-3, 13) is a solution of y = 4 -3x

--------------------------------------------------------------------------------------------------------------
If you need more help, email me at jim_thompson5910@hotmail.com

Also, please consider visiting my website: http://www.freewebs.com/jimthompson5910/home.html and making a donation. Thank you

Jim
--------------------------------------------------------------------------------------------------------------